OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (jiteshmeghwal9):

First find the value of 'x' from one of the equations.

hartnn (hartnn):

isolate y from first equation, y=...? substitute in 2nd equation.

OpenStudy (jiteshmeghwal9):

\[2x+y=1\]\[x=\frac{1-y}{2}\]put this value in other equation

OpenStudy (anonymous):

how do u figure out y

hartnn (hartnn):

u don't, u just find y in terms of x from 1st equation 2x+y=1 y=1-2x ok? now put this in 2nd equation. 4x + 2y = −1

hartnn (hartnn):

any ideas how ?